# Learning Task-Agnostic Skill Bases to Uncover Motor Primitives in Animal Behavior

### v1

@@ 5.2.2 Results
To assess model performance, we estimate the task-specific weights $w(t)$ and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$. The recovered reward functions align closely with the ground truth (Fig. 3A, bottom row). In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it. In the home-seeking task, a distinct peak appears at the home cage. In the exploration task, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
@@
We further visualize the learned skills by applying PCA to obtain PC skills. See the Appendix. C for raw features of skills. Fig. 3B displays two top-contributing skills for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig. 3C). The water-related skills promote movement toward the water port, while the home-related skills guide navigation to the cage. Notably, the skills important for one task have minimal or negative contributions to the other (Fig. 3C), indicating clear functional specialization.
These results show that our model accurately recovers multiple reward functions from real behavior and, importantly, learns interpretable skills whose contributions to each task are both distinct and behaviorally meaningful.
@@ 5.3 Application to animal free-moving behavior
5.3.1 Dataset and model setup
@@
In this section, we apply our method to a dataset of free-moving mouse behaviors to extract motor skills and analyze pose dynamics, specifically how complex pose sequences are dynamically composed of underlying motor primitives. The dataset [2] contains the keypoint coordinates of eight mouse body parts, including four points along the spine, the head, the nose, and both ears. Each dimension of the state corresponds to either $x$ or $y$ coordinate of a body part. The action at each timestep is defined as the velocity of the state, computed as $a_{t}=(s_{t+1}-s_{t})/\delta t$. We set the time interval to $\delta t=1$ ($1/30s$ in the original dataset). We formulate the data using a continuous MDP.
@@
As outlined in Sec. 3.2, we parameterize the transition function as $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})\right)$, and use NCE to learn the feature functions $\psi(s,a)$ and $\nu(s^{\prime})$. We then apply a transformation function $f$ and again use NCE to learn the skill representation $\phi(s,a)=f(\psi(s,a))$ and the time-dependent weights $u(t)$. To ensure temporal smoothness in the learned weights $u(t)$, we place a Gaussian random walk prior over the trajectories: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$. We optimize $f$ and $u(t)$ using coordinate descent, updating them alternately. For this dataset, our focus is on understanding the learned policy structure, which reflects the dynamics of animal poses. Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned skill representation $\phi$ and the temporal weights $u(t)$.
We compare our SKIL method with a baseline model, Keypoint-MoSeq [2], a state-of-the-art behavior segmentation algorithm that partitions continuous pose trajectories into discrete behavioral syllables. For a fair comparison, we train Keypoint-MoSeq on the training set to learn its parameters and then use it to infer behavioral syllables, i.e., perform segmentation, on the test trajectories.
5.3.2 Results
@@
We evaluate performance using the area under the Receiver Operating Characteristic (ROC) curve (AUC), which quantifies the model’s ability to distinguish positive from negative samples. AUC is chosen because it allows direct comparison between our unnormalized energy function and Keypoint-MoSeq’s likelihood score. Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$. We use the log-likelihood of the autoregressive model in Keypoint-MoSeq as its prediction score and the negative energy function $f(\psi(s,a))^{\top}u(t)$ as our prediction score, respectively.
@@
We repeat the experiment 10 times and report the results as a box plot in Fig. 4A. Our model achieves higher AUC scores on both training and test sets (paired t-test, $p_{train}=0.0037,p_{test}=0.0046$), demonstrating a stronger ability to distinguish positive from negative samples. This suggests that SKIL more accurately captures time-varying pose dynamics through smoothly evolving motor primitives, while Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity.
@@
Beyond quantitative comparisons, we also qualitatively visualize and interpret the key skills $\phi(s,a)$ associated with example pose dynamics. We examine the time-varying weights $u(t)$ and compare them with the behavioral syllable segmentation produced by Keypoint-MoSeq (Fig. 4B). We present five example clips of observed animal behavior, each consisting of five consecutive frames. Each clip is characterized by a unique combination of motor skills. For each clip, we show the top 1-2 most dominant skills. For each skill, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that skill. The movement semantics of each skill are labeled above the visualizations. A more comprehensive visualization of all learned skills is included in the Appendix. D.
@@
By combining the discovered skills with real animal behavior, we assess the interpretability of each skill in the five clips (Fig. 4B). First, the right-turn behavior in clip 1 is captured by the dominance of two rightward skills (skill 1 and skill 2). A stronger movement at the head is reflected by a higher value of skill 1. Clips 2, 3, and 4 show the mouse turning right, pausing to groom its head and ears, and then continuing to turn right. The alternate dominance of skill 3 and skill 4 aligns well with the behavior dynamics. Clip 5 shows a simultaneous behavioral mixture of moving forward and sniffing, and is captured by the equal strength of skill 5 and skill 6. Across all skills, skill 4 appears across clips 2, 3, and 4, showing its general utility. The transitions and mixtures of behaviors are effectively reflected in the learned skills and their temporal weights $u(t)$.
@@
However, when we examine the segmentation produced by Keypoint-MoSeq and the syllable dynamics (Fig. 4C), we find inconsistencies. For example, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned. Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements. For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
@@
Taken together, these results show that compared with similar approaches, SKIL provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor skills. This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor primitives.
Figure 4: A. AUC on training and test sets. B. We present example learned skills, behavior clips, $u(t)$, and segmentation from Keypoint-MoSeq. C. Syllables identified by Keypoint-MoSeq (adopted from [2]).
@@ 6 Discussion
We introduce the SKIL framework that advances animal behavior modeling by offering a flexible, interpretable alternative to traditional discrete behavioral segmentation methods. It learns task-agnostic motor primitives (skill sets) that dynamically compose into complex behaviors. SKIL effectively recovers meaningful skills, reconstructs reward functions, and captures nuanced behavioral dynamics, providing a principled and quantitative model of how complex policies arise from fundamental motor components.
Several limitations remain to be addressed in future work. First, the accuracy of inferred skills is sensitive to input data quality, as occlusions or tracking errors can degrade performance. Additionally, training the energy-based models can also be computationally intensive, potentially restricting real-time or large-scale applications. Lastly, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned "skills" and specific neural circuit dynamics still requires further experimental validation.
Beyond advancing animal behavior research, SKIL has broader implications. Positively, a better understanding of motor control mechanisms could, for instance, inform new treatments for movement disorders or inspire more adaptable AI. On the other side, extending these principles to model human behavior carries ethical risks, such as perpetuating or amplifying societal biases present in training data. A robust ethical framework is essential to mitigate such risks in the development and application of these technologies.
Appendix A Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
@@
The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature [6, 35] and obtain the spectral decomposition of Eq. (7) as
@@
Note that Eq. (11) needs infinite $\omega$ to calculate the expectation. To connect it to finite dimension $\phi(s,a)\in\mathbb{R}^{d},\mu(s^{\prime})\in\mathbb{R}^{d}$, we use the Monte-Carlo method to approximate it with finite samples,
@@
Hence, Eq. (7) can, in principle, yield the skill representation introduced earlier.
Appendix B Learned Skills in the Multi-task Gridworld Dataset
Figure 5: A. State-action maps for all 64 skills. B. Reward weight $w$ for all 64 skills.
@@ Appendix B Learned Skills in the Multi-task Gridworld Dataset
We visualize all original 64 skills (Fig. 5) introduced in Sec. 5.1. It shows some meaningful patterns as mentioned before. For example, skill 0 assigns high values to those $(s,a)$ pairs leading to the middle-middle grid and the bottom-middle grid, and assigns low values to the up-left grid and the up-right grid. Thus, it is employed negatively in Task 0 (up-left reward) and Task 2 (up-right reward). However, Task 4 (middle-middle) didn’t use this skill and used skill 39 negatively instead. The complex many-to-many relationship between skills and tasks informs us of the redundancy in the original skills, which inspires us to use PCA to analyze the principal components of the skill space and simplify the skill weights. It could be seen from the comparison between Fig. 2D and Fig. 5A that principal components are a less redundant description of the skill space.
Appendix C Learned Skills in the Animal Navigation Behavior Dataset
Figure 6: A. Left. The value for each $(s,a)$ pair in skill 0 before PCA. Right. The most important $(s,a)$ pairs. B. Left. The value for each $(s,a)$ pair in skill 0 after PCA. Right. The most important $(s,a)$ pairs. C. Boxplot for the effective dimensions before and after PCA.
@@ Appendix C Learned Skills in the Animal Navigation Behavior Dataset
In the original skills of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the skill. Given the redundancy of the skill sets, as in Appendix. B, we perform PCA to analyze the principle components of the skill space and simplify the skill representation. To show the effect of PCA, we plot one skill (skill 0) before (Fig. 6A right) and after PCA (Fig. 6B right). Basically, we only want to show the most important pairs in one map and do not want low-value pairs to disturb the visualization. To determine how many $(s,a)$ pairs are important, we sort the $(s,a)$ pairs based on the value $\phi(s,a)$ in skill 0, i.e., the first feature of the output of $\phi(s,a)$ (Fig. 6A left and B left). It could be seen straightforwardly that after PCA, the skill becomes more concentrated on several $(s,a)$ pairs. We calculate the mean $\mu$ and variance $\sigma$ across all skills and all dimensions and take $\mu+\sigma$ as the threshold, above which $(s,a)$ pairs are deemed the most important ones and are shown on the right. We show 80 pairs before PCA and 8 pairs after PCA. The number of the most important pairs in each skill is called the “effective dimension.” The effective dimension is calculated across all skills (Fig. 6C). Paired t-test ($p=1.3\times 10^{-51}$) shows that there exists a significant decrease of effective dimensions after PCA. So the map becomes more distinct and functionally separated.
Appendix D Learned Skills in the Animal Free-moving Behavior Dataset
Figure 7: A. Policy weight $u(t)$ for all 64 skills. B. The motion field for all 64 skills learned from the video, computed by averaging the states and actions that most strongly activate each skill.

### v2

@@ Results.
To assess model performance, we estimate the task-specific weights $w(t)$ and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$. The recovered reward functions align closely with the ground truth (Fig. 3A, bottom row). In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it. During home-seeking, a distinct peak appears at the home cage. When exploring, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
@@
We further visualize the learned motifs by applying PCA to obtain PC motifs. See the Appendix. E for raw features of motifs. Fig. 3B displays two top-contributing motifs for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig. 3C). The water-related motifs promote movement toward the water port, while the home-related motifs guide navigation to the cage. Notably, the motifs important for one task have minimal or negative contributions to the other (Fig. 3C), indicating clear functional specialization. Fig. 3C also shows non-task-specific motifs with similar weights in both water- and home-seeking tasks.
These results show that our model not only recovers multiple reward functions from real behavior but also learns interpretable motifs whose contributions to each task are distinct and behaviorally meaningful. Unlike previous reward discovery on this dataset (Ashwood et al., 2022; Ke et al., 2025), which cannot identify such motifs, our approach reveals how reward maps can be decomposed into smaller, reusable action components. This decomposition offers a mechanistic view of how local decision processes combine to produce strategies such as water-seeking or home-seeking.
@@ 4.3 Application to animal free-moving behavior
Dataset and model setup.
@@
We finally apply our method to a continuous dataset of free-moving mouse behaviors (Weinreb et al., 2024) to extract motor motifs and analyze fine-grained pose dynamics. This dataset contains the keypoint coordinates of eight mouse body parts, including the head, the nose, both ears, and four spine nodes. Each dimension of the state corresponds to either $x$ or $y$ coordinate of a body part. The action at each timestep is defined as the velocity of state, $a_{t}=(s_{t+1}-s_{t})/\delta t$. We set the time interval to $\delta t=1$ ($1/30s$ in the original dataset). We formulate the data using a continuous MDP. Studying this dataset allows us to ask whether free-moving behaviors, which often appear as mixtures of grooming, locomotion, and postural adjustments, can be represented as combinations of a compact set of task-agnostic motor motifs.
@@
As outlined in Sec. 3.2, we use NCE to learn the motif representation $\phi(s,a)$ and the time-dependent weights $u(t)$. To ensure temporal smoothness in the learned weights $u(t)$, we place a Gaussian random walk prior over the trajectories: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$. To perform more stable optimization, we optimize $\phi(s,a)$ and $u(t)$ using coordinate descent, updating them alternately. For this dataset, our focus is on understanding the learned policy structure, which reflects the dynamics of animal poses. Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned motif representation $\phi$ and the temporal weights $u(t)$. With respect to the choice of the number of motifs, we find in practice that the performance grows more slowly once past $d=64$, so we choose this as an optimal number. See Appendix. F for more training details.
We compare our MCD method with two representative behavior segmentation approaches: (1) Keypoint-MoSeq (Weinreb et al., 2024), as a representative for switching-dynamics-based segmentation methods; and (2) SemiSeg (Whiteway et al., 2021), as a representative for clustering-based behavior segmentation methods. Note that Keypoint-MoSeq is regarded as a SOTA approach, because it extends the autoregressive hidden Markov model (AR-HMM) and MoSeq (Wiltschko et al., 2015), and has been shown to outperform B-SOiD (Hsu & Yttri, 2021), VAME (Luxem et al., 2022), and MotionMapper (Berman et al., 2014). We also include OPAL (Ajay et al., 2021), a representative autoencoder-based motif learning algorithm from robotics, as a baseline to highlight the advantages of our method and why robotics approaches are ill-suited for behavioral segmentation in neuroscience.
Results.
@@
We evaluate performance using the area under the Receiver Operating Characteristic (ROC) curve (AUC), which quantifies the model’s ability to distinguish positive from negative samples. AUC is chosen because it allows direct comparison between our unnormalized energy function and Keypoint-MoSeq’s likelihood score. Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$. With respect to the choice of the prediction score, for Keypoint-Moseq, we use the action log-likelihood. For MCD, we use the negative energy function $\phi(s,a)^{\top}u(t)$. For SemiSeg, we assume action variance=1 and use the Gaussian log-likelihood of actions. For OPAL, we use the action log-likelihood.
@@
We repeat the experiment 10 times and report the results as box plots in Fig. 4A. Our model achieves the highest AUC on both training and test sets (paired t-test, $p<0.05$ for every baseline), demonstrating the strongest ability to distinguish positive from negative samples. This suggests that MCD accurately captures time-varying pose dynamics through smoothly evolving motifs, while other models fail. Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity, and the autoencoders in the other two models have weaker expressive ability.
@@
Beyond quantitative comparisons, we also qualitatively visualize and interpret the key motifs $\phi(s,a)$ associated with example pose dynamics. For MCD, we examine the time-varying weights $u(t)$ of a long animal behavior video (length=250) and choose five example animal behavior clips (length=5) to check the interpretability of the result (Fig. 4B). For our model, each clip is characterized by a unique combination of motor motifs. For each clip, we show the top 1-2 most dominant motifs. For each motif, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that motif. The movement semantics of each motif are labeled above the visualizations. A more comprehensive visualization of all learned motifs is included in the Appendix. F. For comparison, we run other models on the same behavior video, showing the latent representations by SemiSeg (Fig. 4C) and OPAL (Fig. 4D). To show clearer results for SemiSeg and OPAL, we further run KMeans (k=10) on the first 10 PCs of the latents throughout the video and show the segmentation result at the bottom of the latent representations (Fig. 4C, D). We also show the behavioral syllable segmentation produced by Keypoint-MoSeq in Fig. 4E.
@@
By combining the discovered motifs with real animal behavior, we assess the interpretability of each motif in the five clips (Fig. 4B). First, the right-turn behavior in clip 1 is captured by the dominance of two rightward motifs (motif 1 and motif 2). A stronger movement at the head is reflected by a higher value of motif 1. Clips 2, 3, and 4 show the mouse turning right, pausing to groom its head and ears, and then continuing to turn right. The alternate dominance of motif 3 and motif 4 aligns well with the behavior dynamics. Clip 5 shows a simultaneous behavioral mixture of moving forward and sniffing, and is captured by the equal strength of motif 5 and motif 6. Across all motifs, motif 4 appears across clips 2, 3, and 4, showing its general utility. The transitions and mixtures of behaviors are effectively reflected in the learned motifs and their temporal weights $u(t)$.
@@
However, when we examine the segmentation and latent produced by other methods (Fig. 4C-F), we find inconsistencies. In Fig. 4C, alternate dominant behavior patterns in Clip 2-4 in the video are not reflected in its motif weights during this period, as the only dominant motif is the yellow-green one. In Fig. 4D, the motif weights are too dense to interpret. The segmentation results at the bottom of Fig. 4 C and D do not even have repeated behavior patterns and thus could not be seen as a reasonable behavior motif representation. For Keypoint-Moseq, in Fig. 4E, F, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned. Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements. For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
@@
Taken together, these results show that compared with similar approaches, MCD provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor motifs. This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor motifs.
Figure 4: A. AUC on training and test sets. We take an example behavior video and run our algorithm. We visualize the motif weights $u(t)$ and show the representative motifs in B. For the baseline SemiSeg, we show the latent skills and segmentation results in C. For the baseline OPAL, we show the latent skills and segmentation results in D. Then we show the segmentation results of Keypoint-MoSeq (Weinreb et al., 2024) in E and the representative motifs/syllables in F.
@@ 5 Discussion
Several limitations remain to be addressed in future work. First, the accuracy of inferred motifs is sensitive to input data quality, as occlusions or tracking errors can degrade performance. Additionally, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned "motifs" and specific neural dynamics still requires further experimental validation.
Ethics statement
Beyond advancing animal behavior research, MCD has broader implications. Positively, a better understanding of motor control mechanisms could, for instance, inform new treatments for movement disorders or inspire more adaptable AI. On the other side, extending these principles to model human behavior carries ethical risks, such as perpetuating or amplifying societal biases present in training data. A robust ethical framework is essential to mitigate such risks in the development and application of these technologies.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. All source code for model training and evaluation is included in the supplementary material, allowing independent verification and replication of our experiments. The complete set of hyperparameter values is documented in the appendix. Additionally, the data preprocessing procedures and evaluation protocols are described in the main text. These resources provide sufficient information for reproducing the results reported in this paper.
Appendix A LLM Usage
In preparing this manuscript, we employed a large language model (OpenAI ChatGPT, GPT-5) as a writing assistant. The model was used exclusively for polishing English grammar, improving clarity, and suggesting more natural phrasing in certain sections of the text. All scientific content, experimental design, analyses, and interpretations were conceived, written, and verified by the authors. The LLM was not used to generate original research ideas, analyses, or results. To ensure accuracy, all model-suggested edits were carefully reviewed and, where necessary, modified by the authors.
Appendix B Hyperparameter Setting
We train MCD using the following hyperparameters:
General hyperparameters.
  • Discount factor: $\gamma=0.99$
  • Number of epochs: $1\times 10^{6}$
  • Batch size: 256256
Motif representations.
The motif representation $\phi(s,a)\in\mathbb{R}^{d}$ and $\mu(s^{\prime})\in\mathbb{R}^{d}$ were adopted with different motif dimensions $d$ depending on the task:
  • Gridworld: $d=64$
  • Animal navigation: $d=128$
  • Animal free-moving: $d=64$
Model architectures.
  • Discrete version: $\phi$ and $\mu$ are parameterized by one-hidden-layer neural networks (hidden size =512=512).
  • Continuous version:
    • $\phi$: no hidden layer
    • $\mu$: one hidden layer (hidden size =512=512)
    • $\psi$: no hidden layer
Learning rates.
  • Discrete version:
    $\phi:1\times 10^{-3},\quad\mu:1\times 10^{-3},\quad u:3\times 10^{-4},\quad w:3\times 10^{-4}.$
  • Continuous version:
    $\psi:5\times 10^{-4},\quad\nu:5\times 10^{-4},\quad f:3\times 10^{-4},\quad u:3\times 10^{-4}.$
    During testing, $u$ and $f$ are further optimized on the new sequence using gradient descent with learning rate $1\times 10^{-3}$.
We train SemiSeg and OPAL using the following hyperparameters:
  • Discount factor: $\gamma=0.99$
  • Number of epochs: $1\times 10^{6}$
  • Batch size: $4\times 250$ (4 sequences, each of length=250 because this is an RNN-based inference model)
  • Latent dimension: $d=64$
  • Learning rate: $1\times 10^{-4}$ (tuned to get better results)
Besides, the following loss coefficients are shared across three models for interpretability results.
  • Temporal smoothness Gaussian random walk loss: 1010
  • Sparsity L1-loss: $0.1$
Appendix C Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
@@
The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature (Dai et al., 2014; Rahimi & Recht, 2007) and obtain the spectral decomposition of Eq. 6 as
@@
Note that Eq. 10 needs infinite $\omega$ to calculate the expectation. To connect it to finite dimension $\phi(s,a)\in\mathbb{R}^{d},\mu(s^{\prime})\in\mathbb{R}^{d}$, we use the Monte-Carlo method to approximate it with finite samples,
@@
Hence, Eq. 6 can, in principle, yield the motif representation introduced earlier.
Appendix D Multi-task Gridworld Dataset
D.1 Dataset
Figure 5: A. State-action maps for all 64 motifs. B. Reward weight $w$ for all 64 motifs.
D.2 Learned Motifs
@@
We visualize all original 64 motifs (Fig. 5) introduced in Sec. 4.1. It shows some meaningful patterns as mentioned before. For example, motif 0 assigns high values to those $(s,a)$ pairs leading to the middle-middle grid and the bottom-middle grid, and assigns low values to the up-left grid and the up-right grid. Thus, it is employed negatively in Task 0 (up-left reward) and Task 2 (up-right reward). However, Task 4 (middle-middle) didn’t use this motif and used motif 39 negatively instead. The complex many-to-many relationship between motifs and tasks informs us of the redundancy in the original motifs, which inspires us to use PCA to analyze the principal components of the motif space and simplify the motif weights. It could be seen from the comparison between Fig. 2D and Fig. 5A that principal components are a less redundant description of the motif space.
Appendix E Animal Navigation Behavior Dataset
E.1 Learned Motifs
@@
In the original motifs of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the motif. Given the redundancy of the motif sets, as in Appendix. D, we perform PCA to analyze the principle components of the motif space and simplify the motif representation. To show the effect of PCA, we plot one motif (motif 0) before (Fig. 6A right) and after PCA (Fig. 6B right). Basically, we only want to show the most important pairs in one map and do not want low-value pairs to disturb the visualization. To determine how many $(s,a)$ pairs are important, we sort the $(s,a)$ pairs based on the value $\phi(s,a)$ in motif 0, i.e., the first feature of the output of $\phi(s,a)$ (Fig. 6A left and B left). It could be seen straightforwardly that after PCA, the motif becomes more concentrated on several $(s,a)$ pairs. We calculate the mean $\mu$ and variance $\sigma$ across all motifs and all dimensions and take $\mu+\sigma$ as the threshold, above which $(s,a)$ pairs are deemed the most important ones and are shown on the right. We show 80 pairs before PCA and 8 pairs after PCA. The number of the most important pairs in each motif is called the “effective dimension.” The effective dimension is calculated across all motifs (Fig. 6C). Paired t-test ($p=1.3\times 10^{-51}$) shows that there exists a significant decrease of effective dimensions after PCA. So the map becomes more distinct and functionally separated.
Figure 6: A. Left. The value for each $(s,a)$ pair in motif 0 before PCA. Right. The most important $(s,a)$ pairs. B. Left. The value for each $(s,a)$ pair in motif 0 after PCA. Right. The most important $(s,a)$ pairs. C. Boxplot for the effective dimensions before and after PCA.
Appendix F Animal Free-moving Behavior Dataset
F.1 Dataset and training
F.2 Learned Motifs